If A=2xz^2i-yzj+3xz^3k then what is the value of curlA?

1
Expert's answer
2022-01-21T14:31:51-0500

We have

"A=2xz^{2}i-yzj+3xz^{3}k"

To get the value for Curl "A"

We shall find the cross product "\nabla" X "A"


So, we have

"\nabla" X "A" "\displaystyle =\begin{vmatrix}
 i & j & k \\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z} \\
2xz^{2} & -yzj & 3xz^{3}
\end{vmatrix}"


"\displaystyle\\=i\left\{\frac{\partial}{\partial y}(3xz^{3})-\frac{\partial}{\partial z}(-yz)\right\}-j\left\{\frac{\partial}{\partial x}(3xz^{3})-\frac{\partial}{\partial z}(2xz^{2})\right\}+k\left\{\frac{\partial}{\partial x}(-yz)-\frac{\partial}{\partial y}(2xz^{2})\right\}"


"=i\left\{-(-y)\right\}-j\left\{3z^{3}-4xz\right\}+k\left\{0-0\right\}"



"=yi-(3z^{3}-4xz)j+0k\\=yi-z(3z^{2}-4x)j"


Which is the value of Curl A.